OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x3y4)(2x2y6)

OpenStudy (anonymous):

3^2*2^2*y^10

OpenStudy (anonymous):

or 36*y^10

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

u welcome dear

OpenStudy (mayankdevnani):

(3x3y4)(2x2y6) 9y^4*4y^6 36y^10 first we multiply coeficients i.e (3*3) and (2*2)=9*4=36 then (y^4) and (y^6) is added after the multiplication of coefficients then we multiply y^4 and Y^6 i.e y^4+6=y^10(Law of Exponents) so join all these multiplications we get 36y^10

OpenStudy (mayankdevnani):

got it @April18

OpenStudy (anonymous):

yes thank you

OpenStudy (mayankdevnani):

welcome